OpenStudy (anonymous):

What is the probability if I roll an even number less than 6 on a dice

OpenStudy (anonymous):

So it is even number less than 6. so it will be 2 and 4 out of 6, so 2/6 or 1/3 1/3

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Welcome!!! :D and anytime

OpenStudy (anonymous):

I need help again

OpenStudy (anonymous):

what is the probability of a number that is less than 3 on a dice

geerky42 (geerky42):

desired outcomes are 1 and 2, out of 6 total possible outcomes. So that would be \(\dfrac{\operatorname{desired outcomes}}{\operatorname{total outcomes}} = \dfrac{2}{6} = \boxed{\dfrac{1}{3}}\)